find the coordinates of the center of the circles x^2+y^2-2x-4y-31=0
Answer by Cromlix(4381)

The general equation of a circle
= x^2 + y^2 + 2gx + 2fy + c = 0
Centre = (-g,-f)
Therefore the equation of the circle:
x^2 + y^2 - 2x - 4y - 31 = 0
has (1,2) as the coordinates of its centre.
Hope this helps.
:-)